Solve the following differential equations: (i) (𝑦 3 + 1) 𝑑𝑦 𝑑π‘₯ βˆ’ π‘₯𝑦 2 = π‘₯, (ii) π‘₯𝑦 2 𝑑𝑦 𝑑π‘₯ = π‘₯ 3 βˆ’ 𝑦 3, (iii) (1 + sin π‘₯) 𝑑𝑦 𝑑π‘₯ + 𝑦 cos π‘₯ = tan π‘₯, (iv) π‘₯ 𝑑𝑦 𝑑π‘₯ βˆ’ 5𝑦 = π‘₯ 7.


1
Expert's answer
2021-10-22T11:06:52-0400

(i)

"(y^{3}+1)\frac{dy}{dx}=x+xy^{2}"

"(y^{3}+1)\frac{dy}{dx}=x(1+y^{2})"

"\frac{1+y^{3}}{1+y^{2}}dy=xdx"

Integrate on both sides

"\int \frac{1+y^{3}}{1+y^{2}}dy=\int xdx"

Let "1+y^{2}" =t

2ydy=dt

ydy="\frac{dt}{2}"

"y^{2}=(t-1)"

"\int \frac{1}{1+y^{2}}dy+\int \frac{y^{3}}{1+y^{2}}=\int xdx"

"tan^{-1}(y)+\frac{1}{2}\int \frac{(t-1)}{t}dt=\frac{x^{2}}{2}+C"

"tan^{-1}(y)+\frac{1}{2}\int (1-\frac{1}{t})dt=\frac{x^{2}}{2}+C"

"tan^{-1}(y)+\frac{1}{2}(t-ln\ t)=\frac{x^{2}}{2}+C"

Putting the values of t answer becomes;

"tan^{-1}(y)+\frac{1}{2}[(1+y^{2})-ln(1+y^{2})]=\frac{x^{2}}{2}+C"


(ii)

"xy^{2}\frac{dy}{dx}=x^{3}-y^{3}"

"\frac{dy}{dx}=\frac{x^{3}-y^{3}}{xy^{2}}"

On dividing the numerator and denominator by x3

"\frac{dy}{dx}=\frac{1-(\frac{y}{x})^{3}}{(\frac{y}{x})^{2}}"

Let "\frac{y}{x}=V" "\implies y=Vx"

"\frac{dy}{dx}=V+x\frac{dV}{dx}"

"V+x\frac{dV}{dx}=\frac{1-V^{3}}{V^{2}}"

"x\frac{dV}{dx}=\frac{1-V^{3}}{V^{2}}-V"

"x\frac{dV}{dx}=\frac{1-V^{3}-V^{3}}{V^{2}}"

"x\frac{dV}{dx}=\frac{1-2V^{3}}{V^{2}}"

"\frac{V^{2}}{1-2V^{3}}dV=\frac{1}{x}dx"

Integrating both sides

"\int \frac{V^{2}}{1-2V^{3}}dV=\int \frac{1}{x}dx"

Let 1-2V3=t

-6V2dV=dt

"V^{2}dV=-\frac{1}{6}dt"

"\frac{-1}{6}\int \frac{1}{t}dt=\int \frac{1}{x}dx"

"-\frac{1}{6}ln\ t=ln\ x+C"

"-\frac{1}{6}ln\ (1-2V^{3})=ln\ x+C"

"-\frac{1}{6}ln\ (1-2(\frac{y}{x})^{3})=ln\ x+C"

"-\frac{1}{6}ln[\frac{x^{3}-2y^{3}}{x^{3}}]=ln\ x+C"


(iii)

"(1+sin\ x)\frac{dy}{dx}+ycos\ x=tan\ x"

"\frac{dy}{dx}+\frac{ycos\ x}{(1+sin\ x)}=\frac{tan\ x}{(1+sin\ x)}"

This is a linear equation in y

Integrating factor ,I.F="e^{\int{\frac{cos\ x}{1+sin\ x}}dx}"

"=e^{ln(1+sin\ x)}"

"=1+sin\ x"

Multiply through by I.F to get;

"\frac{d}{dx}[y(1+sin\ x)]=\frac{tan\ x}{1+sin\ x}(1+sin\ x)"

"\frac{d}{dx}[y(1+sin\ x)]={tan\ x}"

Integrating;

"y(1+sin\ x)=\int tan\ xdx"

"y(1+sin\ x)=ln|sec\ x|+C"

Our solution therefore is;

"y=\frac{ln|sex\ x|}{1+sin\ x}+\frac{C}{1+sin\ x}"

(iv)

"x\frac{dy}{dx}-5y=x^{7}"

"\frac{dy}{dx}-\frac{5y}{x}=x^{6}"

This is in the form "\frac{dy}{dx}+Py=Q"

"P=\frac{-5}{x},Q=x^{6}"

Integrating factor, I.F="e^{\int Pdx}"

I.f="e^{-5{\int \frac{1}{x}dx}}=e^{-5ln\ x}=e^{ln\ x^{-5}}=\frac{1}{x^{5}}"

"y*I.f=\int QI.f dx"

"y*\frac{1}{x^{5}}=\int x^{6}*\frac{1}{x^{5}}dx"

"\frac{y}{x^{5}}=\int xdx"

"\frac{y}{x^{5}}=\frac{x^{2}}{2}+C"

"y=\frac{x^{7}}{2}+Cx^{5}"